OpenStudy (mathmath333):

if \(\Large x\geq y\) and \(\Large y>1\) , then value of the expression \(\Large\tt \begin{align} \color{black}{\log_x \dfrac{x}{y} +\log_y \dfrac{y}{x} }\end{align}\) can never be \(a. -1\\ b. 0.5\\ c. 0\\ d. 1\)

OpenStudy (mathmath333):

i got this \(\Large\tt \begin{align} \color{black}{\log_x \dfrac{x}{y} +\log_y \dfrac{y}{x} \\ =\log_x (x)- \log_x (y)+\log_y (y)+\log_y (x)\\ =2- \log_x (y)-\log_y (x)\\ =2- \dfrac{\log_y (y)}{\log_y (x)}-\log_y (x)\\ =2- \dfrac{1}{\log_y (x)}-\log_y (x)\\ =2- \dfrac{1}{\log_y (x)}-\dfrac{(\log_y (x))^2}{\log_y (x)}\\ =2- \dfrac{1}{\log_y (x)}-\dfrac{2\log_y (x)}{\log_y (x)}\\ =-\dfrac{1}{\log_y (x)} }\end{align}\)

OpenStudy (mathmath333):

@campbell_st

OpenStudy (mathmath333):

is my method wrong

OpenStudy (michele_laino):

After some steps, I found the equivalent form of your expression, here is: your expression = \[-\frac{ (\log _{x}y-1)^{2} }{ \log _{x}y}\], from which, I excluded the values 0.5 and 1

OpenStudy (michele_laino):

that's right?

OpenStudy (mathmath333):

yes thats right but u swapped x and y

OpenStudy (michele_laino):

I have changed the base of the logarithm at the second term, from x to y!

OpenStudy (mathmath333):

ok but i dont get how u excluded options 0.5 and 1

OpenStudy (michele_laino):

In my expression, please change variable in this way: \[\log _{x}y=k\] being k the new variable, now write my expression as a function of k, namely: \[-\frac{ (k-1)^{2} }{ k }\] finally, substitute to my expression your values, you will find an quadratic equation in k, and in some cases there is no solutiopn for k.

OpenStudy (michele_laino):

Sorry I wanted say, equate my expression to your values...

OpenStudy (mathmath333):

m still confused

OpenStudy (loser66):

the 3rd lines copied wrong

OpenStudy (loser66):

it's 2-log_x (y)+ log_y (x) , you did - for the last term

OpenStudy (michele_laino):

for example, for value -1, we can write: \[-\frac{ (k-1)^{2} }{ k }=-1\] now try to solve that equation for k, please

OpenStudy (loser66):

however, I don't know why you have to do a long argument like that,

OpenStudy (loser66):

log > 0 for all x, y the first one is >0, the second one is >0 also, how the sum <0 (I mean option a)

OpenStudy (loser66):

get me? positive number + positive number CANNOT BE negative one then, the so-called "Never be" is a)

OpenStudy (loser66):

no?? hihihi....

OpenStudy (mathmath333):

is this correct \(\Large\tt \begin{align} \color{black}{\log_x \dfrac{x}{y} +\log_y \dfrac{y}{x} \\ =\log_x (x)- \log_x (y)+\log_y (y)-\log_y (x)\\ =2- \log_x (y)-\log_y (x)\\ =2- \dfrac{\log_y (y)}{\log_y (x)}+\log_y (x)\\ =2- \dfrac{1}{\log_y (x)}+\log_y (x)\\ =2- \dfrac{1}{\log_y (x)}+\dfrac{(\log_y (x))^2}{\log_y (x)}\\ =2- \dfrac{1}{\log_y (x)}+2\dfrac{\log_y (x)}{\log_y (x)}\\ =-\dfrac{1}{\log_y (x)} }\end{align}\)

OpenStudy (loser66):

nope

OpenStudy (mathmath333):

why :|

OpenStudy (mathmath333):

i think the second line was wrong third was correct

OpenStudy (mathmath333):

btw what is min and max value for any log

OpenStudy (loser66):

\(log _x y^2 = 2log_x y\) but \((log_x y)^2 \neq 2log_x y\) for example \(log_{10} 100 =2 \), that is \(\log_{10}10^2=2log_{10}10\)

OpenStudy (loser66):

but \((log_{10} 10)^2 =1\)

OpenStudy (loser66):

so that, you can't take down 2 on the 6th line

OpenStudy (mathmath333):

@Michele_Laino \(k ~will ~be+~\)\)

OpenStudy (michele_laino):

the values of k are not important, what is important is if the equation admits or no, solutions for k

ganeshie8 (ganeshie8):

\[\large \rm \log (x^2) = 2\log(x) \ne \left[\log(x)\right]^2\]

OpenStudy (mathmath333):

ok i got it :)

ganeshie8 (ganeshie8):

\[\large \rm \left[\log(x)\right]^2 = \log(x) \times \log(x)\]

ganeshie8 (ganeshie8):

\[\large \rm \log(x^2) = \log(x\cdot x) = \log(x) + \log(x)\]